# Armijo Line-search Makes (Stochastic) Gradient Descent Go Fast

### v1

@@ Appendix E Proofs for Section 6
$\leq f(u)+\frac{L}{2}\left[\left\|u\right\|_{2}^{2}-C\,\mathbb{E}\sum_{t=0}^{T-1}\left[\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right]+\frac{2\,C\,f(u)\,T}{\epsilon}\right]$
Using that $f(u)\leq\epsilon^{2}$ and $\left\|u\right\|=\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon^{2}}\right)$,
@@
$\implies\mathbb{E}[f(\theta_{T})]$ $\leq\epsilon^{2}+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}-C\,\mathbb{E}\sum_{t=0}^{T-1}\left[\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right]+2\,C\,\epsilon\,T\right]$
We also know that for all $t\leq T$, $f(\theta_{t})-f(u)\geq\epsilon$, meaning that $\mathbb{E}\left[\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right]\geq\frac{\epsilon}{f(u)+\epsilon}$. Using this relation,
$\leq\epsilon^{2}+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}-C\,T\,\left[\frac{\epsilon}{f(u)+\epsilon}\right]+2\,C\,\epsilon\,T\right]$
$\leq\epsilon^{2}+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}-\frac{C\,T}{2}+2\,C\,\epsilon\,T\right]$ (Since $f(u)\leq\epsilon^{2}\leq\epsilon$)
@@
$\mathbb{E}[f(\theta_{T})]$ $\leq\epsilon^{2}+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}-\frac{3C\,T}{8}\right]$ (Since $\epsilon\leq\frac{1}{8}$)
@@
Hence, in order to ensure that $f(\theta_{T})\leq\epsilon\,(1+\epsilon)\leq 2\epsilon$, it is sufficient to set $T$ to ensure that,
@@
$\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}-\frac{3C\,T}{8}\right]\leq\epsilon$
Hence, it is sufficient to set $T$ as
@@
$T$ $\geq\frac{8}{3C}\,\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}=\frac{16\,c}{(1-c)\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}$ (where we used $\lambda_{1}=\frac{6}{1-c}$ in $C$)

### v2

@@ C.2 Helper Lemmas
Hence, we need to bound $\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]\right]$ conditioned on the event $G^{c}=\{\inf_{t\in[T]}[f(\theta_{t})-f(u)]\geq\epsilon\}$.
Conditioned on $G^{c}$, we know that for all $t\in[T]$, $f(\theta_{t})-f(u)\geq\epsilon$. Since $f(u)\leq f(\theta_{t})$ for all $t\in[T]$, we can use Lemma 3 with the comparator $u=\beta u^{*}$ and obtain the following inequality for SGD-SLS with $\theta_{0}=0$, where $C=\frac{2c-1}{c}\,\frac{1}{\lambda_{1}}$
$\mathbb{E}[\left\|\theta_{T}-u\right\|_{2}^{2}]$ $\leq\left\|u\right\|_{2}^{2}-\sum_{t=0}^{T-1}\mathbb{E}\left[\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,[f(\theta_{t})-f(u)]\right]+2\eta_{\max}\,\chi^{2}(u)\,T$
$\leq\left\|u\right\|_{2}^{2}-\sum_{t=0}^{T-1}\mathbb{E}\left[\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,[f(\theta_{t})-f(u)]\right]+2\eta_{\max}\,f(u)\,T$ (Since $\chi^{2}(u)\leq f(u)$)
Setting $\eta_{\max}=\frac{C}{\epsilon}$ ensures that, conditioned on $G^{c}$, $\eta_{\max}\geq\frac{C}{f(\theta_{t})}$ for all $t\leq T$. Hence,
$\leq\left\|u\right\|_{2}^{2}-C\,\sum_{t=0}^{T-1}\mathbb{E}\left[\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right]+\frac{2C\,f(u)}{\epsilon}\,T$
Since $f(u)\leq\epsilon^{2}\leq\epsilon$ and conditioned on $G^{c}$, $\left[\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right]\geq\frac{\epsilon}{f(u)+\epsilon}$,
$\leq\left\|u\right\|_{2}^{2}-C\,\sum_{t=0}^{T-1}\mathbb{E}\left[\frac{\epsilon}{f(u)+\epsilon}\right]+\frac{2C\,f(u)}{\epsilon}\,T$
$\leq\beta^{2}-\frac{C\,T}{2}+2C\,\epsilon\,T$ (Since $u=\beta u^{*}$ and $f(u)\leq\epsilon^{2}\leq\epsilon$)
@@
$\implies\mathbb{E}[\left\|\theta_{T}-u\right\|_{2}^{2}]$ $\leq\beta^{2}-\frac{C\,T}{4}$ (For $\epsilon\leq\frac{1}{8}$)
Since $f$ satisfies Assn. 1, 2 and 3, $f$ is $1$ uniform smooth and $f(u)\leq\epsilon$, then, using Lemma 5 with $M=\frac{1}{8}$,
@@
$\mathbb{E}[f(\theta_{T})-f(u)]$ $\leq\frac{\epsilon}{2}+\underbrace{[\nu^{2}+1]}_{:=L}\,\mathbb{E}\frac{\left\|\theta_{T}-u\right\|_{2}^{2}}{2}\leq\frac{\epsilon}{2}+\frac{L}{2}\,\left[\beta^{2}-\frac{C\,T}{4}\right]$
@@
Hence, in order to ensure that $\mathbb{E}[f(\theta_{T})-f(u)]\leq\frac{3\epsilon}{2}$, it is sufficient to set $T$ as follows.
@@
$T$ $\geq\frac{4\,\beta^{2}}{C}=\frac{4}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}\,\frac{c\lambda_{1}}{(2c-1)}$ (Substituting the value of $\beta$ and $C$)
@@
$=\frac{4}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}\,\frac{3\,c\,L_{1}\,(\nu+1)}{(2c-1)\,(1-c)}$ (Substituting the value of $\lambda_{1}$)
$=\frac{864}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}\,\frac{c}{(2c-1)\,(1-c)}$ (Substituting $L_{1}=8$, $\nu=8$)
Hence, with this choice of $T$, we can conclude that,
$\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]|G^{c}\right]$ $\leq\frac{3\epsilon}{2}$
Combining this with Eq. 37, we conclude that
$\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]\right]$ $\leq 3\epsilon\implies\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]\leq 4\,\epsilon$ (Since $f(u)\leq\epsilon$)